# Reverse Probing: Evaluating Knowledge Transfer via Finetuned Task Embeddings for Coreference Resolution

### v1

@@ Abstract
In this work, we reimagine classical probing to evaluate knowledge transfer from simple source to more complex target tasks. Instead of probing frozen representations from a complex source task on diverse simple target probing tasks (as usually done in probing), we explore the effectiveness of embeddings from multiple simple source tasks on a single target task. We select coreference resolution, a linguistically complex problem requiring contextual understanding, as focus target task, and test the usefulness of embeddings from comparably simpler tasks tasks such as paraphrase detection, named entity recognition, and relation extraction. Through systematic experiments, we evaluate the impact of individual and combined task embeddings.
Our findings reveal that task embeddings vary significantly in utility for coreference resolution, with semantic similarity tasks (e.g., paraphrase detection) proving most beneficial. Additionally, representations from intermediate layers of fine-tuned models often outperform those from final layers. Combining embeddings from multiple tasks consistently improves performance, with attention-based aggregation yielding substantial gains. These insights shed light on relationships between task-specific representations and their adaptability to complex downstream tasks, encouraging further exploration of embedding-level task transfer.* Equal contribution.
@@ 1 Introduction
A main intuition behind probing is to evaluate to what degree the representations that are learned from the complex source task can be re-purposed to solve a new, simpler task Belinkov (2022). In our work we decide to reverse this paradigm (reverse probing) and investigate how different source task embeddings, from a model trained on simple tasks, can be adapted to a new, more complex target task. In other words, we try to answer the question: Can we reuse knowledge from simpler tasks for a more complex task? Such a recycling of knowledge is not only interesting to deepen our understanding of what type of knowledge is encoded in language models, but it also results in more energy-efficient deep learning, by reusing network weights and representations.
We choose coreference resolution Lee et al. (2017) as our target task because solving coreference is - up to date - a challenging NLP problem that even newer large language models struggle with Bohnet et al. (2023); artinelli-etal-2024-maverick. Coreference resolution involves understanding of context, what counts as a valid mention and which mentions refer to the same entity. Solving coreference may require different types of linguistic knowledge. Our goal is to find out which types of information from which source task models are useful and how this information can be combined and/or adapted to work for the target task.
@@ 3.2 Source Tasks
We focus on the comparison with standard BERT as well as four other task-specific models.
@@ 4.3 How to Combine Task Embeddings
We employ two different aggregation strategies to combine the embeddings of the source task models: mean and attention-based aggregation. We also experimented with sum instead of mean as an aggregator but the results were comparable and slightly worse than the mean aggregation. E.g., a combination of frozen MRPC with BERT achieves 62.34 F1 with sum and 63.26 with mean (these are average values across three runs with different seeds). Hence, in all further experiments we focus on the comparison between the mean and attention-based aggregation.
@@ 4.4 How to Extract Embeddings
We also consider different ways of embedding manipulations since the final layers of BERT-based models might be too specialized on their corresponding tasks, so that their representations are no longer useful for coreference resolution. In fact, after comparing the embeddings from layer 6 to 12 we found that the best performing layer on our probing task was typically not the final one. E.g., it was layer 9 for MRPC and RE, layer 8 for QA and 6 for NER (see Figure 5). Tables 1, 2 and 3 show the detailed comparisons between the original (full) model outputs as well as the normalized and truncated (to the “best" layer) versions for single models and their combinations (see also Figure 6 and 7 for the plot comparison). Truncation seems to be a good strategy for embedding aggregation and consistently yields best results across different settings. E.g., NER truncated to the “best" layer improves by +26.2% and QA improves by +14% (Figure 6). Also combinations of models work better with truncation, e.g., RE+QA pair gains +8.17% F1 with mean aggregation and +2.84% with attention aggregation when both models are truncated (Figure 7).
@@ 6 Conclusion
We hope that our experiments can help to clarify the impact of the embeddings and their combinations on the target coreference task. We also hope that the “upside-down" probing idea will facilitate further research on finding useful information in the task-specific representations that come from different fine-tuned models.

### v2

@@ Abstract
In this work, we reimagine classical probing to evaluate knowledge transfer from simple source to more complex target tasks. Instead of probing frozen representations from a complex source task on diverse simple target probing tasks (as usually done in probing), we explore the effectiveness of embeddings from multiple simple source tasks on a single target task. We select coreference resolution, a linguistically complex problem requiring contextual understanding, as focus target task, and test the usefulness of embeddings from comparably simpler tasks such as paraphrase detection, named entity recognition, and relation extraction. Through systematic experiments, we evaluate the impact of individual and combined task embeddings.
Our findings reveal that task embeddings vary significantly in utility for coreference resolution, with semantic similarity tasks (e.g., paraphrase detection) proving most beneficial. Additionally, representations from intermediate layers of fine-tuned models often outperform those from final layers. Combining embeddings from multiple tasks consistently improves performance, with attention-based aggregation yielding substantial gains. These insights shed light on relationships between task-specific representations and their adaptability to complex downstream tasks, encouraging further exploration of embedding-level task transfer. Our source code is publicly available.111github.com/Cora4NLP/multi-task-knowledge-transfer * Equal contribution.
@@ 1 Introduction
A main intuition behind probing is to evaluate to what degree the representations that are learned from the complex source task can be re-purposed to solve a new, simpler task Belinkov (2022). In our work we decide to reverse this paradigm (thus reverse probing) and investigate how different source task embeddings, from a model trained on simple tasks, can be adapted to a new, more complex target task. In other words, we try to answer the question: Can we reuse knowledge from simpler tasks for a more complex task? Such a recycling of knowledge is not only interesting to deepen our understanding of what type of knowledge is encoded in language models, but it also results in more energy-efficient deep learning, by reusing network weights and representations.
We choose coreference resolution Lee et al. (2017) as our target task because solving coreference is - up to date - a challenging NLP problem that even newer large language models struggle with Bohnet et al. (2023); Martinelli et al. (2024). Coreference resolution involves understanding of context, what counts as a valid mention and which mentions refer to the same entity. Solving coreference may require different types of linguistic knowledge. Our goal is to find out which types of information from which source task models are useful and how this information can be combined and/or adapted to work for the target task.
@@ 3.2 Source Tasks
We focus on the comparison with standard BERT as well as four other task-specific models. As source tasks we take a range of tasks of varying complexity: Paraphrase identification, named entity recognition, relation extraction, and - a (more complex) source task - quesion answering.
@@ 4.3 How to Combine Task Embeddings
We employ two different aggregation strategies to combine the embeddings of the source task models: mean and attention-based aggregation. Additionally, we experimented with summing instead of using the mean, but the results were comparable or slightly worse: A combination of frozen MRPC with BERT achieves 62.34 F1 with sum and 63.26 with mean (average values across three runs with different seeds). Hence, in all further experiments we focus on the comparison between the mean and attention-based aggregation.
@@ 4.4 How to Extract Embeddings
We also consider different ways of embedding manipulations since the final layers of BERT-based models might be too specialized on their corresponding tasks, so that their representations are no longer useful for coreference resolution. In fact, after comparing the embeddings from layer 6 to 12 we found that the best performing layer on our probing task was typically not the final one. E.g., it was layer 9 for MRPC and RE, layer 8 for QA and 6 for NER (see Figure 5). Tables 1, 2 and 3 show the detailed comparisons between the original (full) model outputs as well as the normalized and truncated (to the “best" layer) versions for single models and their combinations (see also Figure 6 and 7 for the plot comparison). Truncation seems to be a good strategy for embedding aggregation and consistently yields best results across different settings. Truncation improves NER by up to +26.2%. QA is improved by +14% (Figure 6). Also combinations of models work better with truncation, e.g., RE+QA pair gains +8.17% F1 with mean aggregation and +2.84% with attention aggregation when both models are truncated (Figure 7).
@@ 6 Conclusion
In the future it would be interesting to experiment with more types of embedding manipulations. Also, a combination of truncation and normalization could possibly bring some gains for single models. Moreover, it would be interesting to check the effects of attention aggregation with hidden layer concatenations for multiple models (e.g., RE+MRPC). So far we tested the approach that proved to be successful only for single models. Finally, it would be interesting to replicate our experiments on larger (non-BERT) models and tasks (e.g., semantic role labeling, discourse relation classification etc.).
We hope that our experiments can help to clarify the impact of embeddings and their combinations on the target coreference task. We also hope that the reverse probing idea will facilitate further research on finding useful information in the task-specific representations that originate from different fine-tuned models.
Limitations
While this work sheds light on the potential of reverse probing and task embeddings, some limitations arise.
First, we exclusively work with BERT-based models. This gives us a controlled setup, but it also means our findings might not fully translate to largermodels or other architectures like GPT, T5, or multilingual models. Future work needs to investigate a broader range of models.
Our choice of source tasks Paraphrase Detection (MRPC), Named Entity Recognition (NER), Relation Extraction (RE), and Question Answering (QA) - is not exhaustive. There are many other NLP tasks, such as sentiment analysis, syntactic parsing, or commonsense reasoning, that might contribute useful embeddings for coreference resolution. Also, some of the tasks are not necessary simpler tha coreference resolution (e.g., QA), which we pickedas our target task. Generally, our conclusions are centered around coreference resolution as the target task. While this is a challenging and linguistically complex problem, our approach may not directly apply to other NLP tasks with different dependencies, such as machine translation or text summarization. Testing on additional target tasks would be a logical next step.
Lastly, there’s the question of computational efficiency. While we worked with relatively small models, combining embeddings from multiple layers and tasks does introduce extra processing overhead. Finding ways to leverage task embeddings efficiently, without significantly increasing inference costs, is another area worth exploring.
Acknowledgments
This work has been supported by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005).